OpenStudy (anonymous):

Related Rates Problem::Perimeter and Area of a rectangle... The length of a rectangle is increasing at 5 meters per second while the width is decreasing at 3 meters per second. At a certain instant, the length is 2 meters and the width is 10 meters. At that time, determine the rate of change of the area and perimeter.

OpenStudy (anonymous):

DA/dt = W(dW/dt)(L) + L(dL/dt)(W) is this equation right?

OpenStudy (anonymous):

dP/dt = 2W(dW/dt) + 2L(dL/dt) is this right?

OpenStudy (anonymous):

actually, it appears as I was using the chain rule instead of the product rule, I get 44 for the area rate, is this right?

OpenStudy (anonymous):

I get 4 m/s for the perimeter rate...is this right?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

what variables are you using?

OpenStudy (misty1212):

looks like \(W\) and \(L\) so \[A=WL\\ A'=W'L+L'W\] by the product rule

OpenStudy (misty1212):

you are given all the numbers, go ahead and plug them in

OpenStudy (anonymous):

@misty1212 do those answers look correct to you?

OpenStudy (misty1212):

hmm not the first one

OpenStudy (misty1212):

it is not \(WW'L\) just \(W'L\) etc

OpenStudy (misty1212):

it is a product, the product rule says \[A=WL\\ A'=W'L+L'W\]

OpenStudy (anonymous):

sorry, yes, corrected that before plugging in, the answer is adjusted for that formula

OpenStudy (anonymous):

answers are just below there

OpenStudy (misty1212):

oh i didn't plug in the numbers i can check though

OpenStudy (misty1212):

\[L'=5,W'=-3, L-2,W=10\] right ?

OpenStudy (misty1212):

\[-3\times 2+5\times 10=44\]

OpenStudy (misty1212):

\[P=2L+2W\\ P'=2W'+2L'\]

OpenStudy (misty1212):

that one is even easier

OpenStudy (anonymous):

thanks for assisting

OpenStudy (misty1212):

\[2\times5+2\times -3\] right?

OpenStudy (misty1212):

yw \(\color\magenta\heartsuit\)